• What is the reminder? when x4- 3x2- 5x is divided by 2x-1
1
Expert's answer
2021-11-30T04:53:14-0500
"\dfrac{x^4- 3x^2- 5x}{2x-1}"

"=\dfrac{1}{2}x^3+\dfrac{x^4-x^4+\dfrac{1}{2}x^3- 3x^2- 5x}{2x-1}"

"=\dfrac{1}{2}x^3+\dfrac{1}{4}x^2+\dfrac{\dfrac{1}{2}x^3-\dfrac{1}{2}x^3+\dfrac{1}{4}x^2- 3x^2- 5x}{2x-1}"

"=\dfrac{1}{2}x^3+\dfrac{1}{4}x^2-\dfrac{11}{8}x+\dfrac{-\dfrac{11}{4}x^2+\dfrac{11}{4}x^2-\dfrac{11}{8}x-5x}{2x-1}"

"=\dfrac{1}{2}x^3+\dfrac{1}{4}x^2-\dfrac{11}{8}x-\dfrac{51}{16}+\dfrac{-\dfrac{51}{8}x+\dfrac{51}{8}x-\dfrac{51}{16}}{2x-1}"

"=\dfrac{1}{2}x^3+\dfrac{1}{4}x^2-\dfrac{11}{8}x-\dfrac{51}{16}+\dfrac{-\dfrac{51}{16}}{2x-1}"



The reminder when "x^4- 3x^2- 5x" is divided by "2x-1" is "-\dfrac{51}{16}."